You are a day trader on the stock exchange and you have invested all of your life savings in a portfolio containing 6 stocks. After dwelling on the risks involved in your investments, you have come to the conclusion that if at least half of the stock in you portfolio are profitable, your investment wil be a success.
Based on the emprical data available from the historical records of the stock exchange you beleive that the probability of a stock being profitable is 0.39. Assuming that the profitability of each stock is independant of all the other stocks, CALCULATE THE PROBABILITY (S) THAT YOUR INVESTMENT IS A SUCCESS.
THANK YOU

Answer by venugopalramana(3286)

You are a day trader on the stock exchange and you have invested all of your life savings in a portfolio containing 6 stocks. After dwelling on the risks involved in your investments, you have come to the conclusion that if at least half of the stock in you portfolio are profitable, your investment wil be a success.
Based on the emprical data available from the historical records of the stock exchange you beleive that the probability of a stock being profitable is 0.39. Assuming that the profitability of each stock is independant of all the other stocks, CALCULATE THE PROBABILITY (S) THAT YOUR INVESTMENT IS A SUCCESS.
THANK YOU
SO WE NEED ATLEAT 50% OF 6 =6*50/100=3 TO SUCCEED...P SUCESS =0.39..P FAILURE=1- 0.39 = 0.61
P3=(0.39^3){0.61^3)=0.013464
P4=(0.39^4)(0.61^2)=0.008608
P5=(0.39^5)(0.61)=0.005504
P6=(0.39^6)=0.003519
TOTAL P SUCCESS = 0.031095